OpenStudy (anonymous):

The domain of the sine function y = 2sin x is _____. A. 2π B. [-1,1] C. π D. All real numbers

OpenStudy (akshay_budhkar):

all real numbers

OpenStudy (anonymous):

thanks

OpenStudy (akshay_budhkar):

sinx is continuous for all values of x

OpenStudy (anonymous):

domain is from \[(-\frac{\pi}{2},\frac{\pi}{2})\]Principle one

OpenStudy (akshay_budhkar):

ishaan priciple domain is not applicable in this case.. all real numbers is right :)

OpenStudy (anonymous):

yeah but right all real numbers

OpenStudy (akshay_budhkar):

principle domain is used only in formulae..

OpenStudy (anonymous):

write i mean

OpenStudy (akshay_budhkar):

ishaan what are you telling? lol i am not getting u

OpenStudy (anonymous):

well i said right but it should have been write